# An Objective Improvement Approach

### v4

@@ 5.3. Games with Efficient Objective Improvement
The second condition, which allows us to identify better strategies efficiently, is to assume that, for every strategy $\sigma$ for both players, if a solution $\mathsf{val}$ defined by a basis is not optimal for $f_{\sigma}$ under the constraints $H$, then there is a single basis change that improves it. We call such games improving. Being improving is the property we really need, while sharpness is a property that allows us to almost surely make a game improving by applying small noise on the objective functions.
@@
(Improving Game) A game $\mathcal{G}$ is called $\sigma$-improving for an objective function $f_{\sigma}$ defined by a joined strategy $\sigma$ and a non-optimal solution $\mathsf{val}$ defined by a basis for $f_{\sigma}$ if, and only if, there is a single basis change that leads to a better strategy (having a smaller value of $f_{\sigma}$) defined by the new basis.
@@
Let $\mathcal{G}=(V_{\min},V_{\max},E,w,\lambda)$ be an improving DPG, $\sigma$ a strategy for both players that is not co-optimal, $\mathsf{val}$ an optimal solution returned at Line 5 of Algorithm 2 for $f_{\sigma}$, and let there be no local improvements of $\sigma$ for $\mathsf{val}$. Then there is (at least) one neighbouring solution $\mathsf{val}^{\prime\prime}$ to $\mathsf{val}$ such that there is a better strategy $\sigma^{\prime}$ that satisfies $f_{\sigma^{\prime}}(\mathsf{val}^{\prime\prime})<f_{\sigma}(\mathsf{val})$.
This strategy $\sigma^{\prime}$ can be selected in such a way that $(v,\sigma^{\prime}(v))\in E_{\mathsf{val}}^{\sigma}$ holds for all $v\in V$ for a given set $E_{\mathsf{val}}^{\sigma}$.
Proof 5.3.
@@
Remark.
As an alternative to considering all neighbouring bases of $\mathbf{b}$, determining their valuation (and checking that they are solutions), and determining the objective function for $\sigma^{\prime}$, we can also determine the reduced cost for exchanging a basis based on the best improvement obtained by any function derived from $E^{\sigma}_{\mathsf{val}}$ and use this to locate an improvement.
5.4. Making Games Sharp and Improving
Naturally, not every game is sharp or improving. In this subsection, we first discuss how to almost surely make games sharp by adding sufficiently small random noise to the edge weights, and then discuss how to treat games that are not improving by assigning randomly chosen factors, with which the offsets of edges are weighted. Note that these are ‘global’ adjustments of the game that only need to be applied once, as it is the game that becomes sharp and improving, respectively.
Starting with the small noise to add on the edge weights, we first create notation for expressing how much we can change edge weights, such that joint co-optimal strategies of the resulting game are joint co-optimal strategies in the original game. To this end, we define the gap of a game.
@@ 5.4. Making Games Sharp and Improving
This defines a run $\rho^{\prime}=d_{1}d_{2}d_{3}\ldots$ (with $d_{1}=e$), again in the form of a “lasso path” that consists of an initial path $d_{1},\ldots,d_{k^{\prime}}$, followed by an infinitely often repeated cycle $d_{1}^{\prime},\ldots,d_{\ell^{\prime}}^{\prime}$, where the only edge that occurs twice on the path $d_{1},\ldots,d_{k^{\prime}},d_{1}^{\prime},\ldots,d_{\ell^{\prime}}^{\prime}$ is the edge reached before $d_{1}^{\prime}$ and after $d_{\ell^{\prime}}^{\prime}$, while all other edges occur only once.
@@
The two loops are either the same or disjoined, and in both cases they have a common multiplier smaller than $\prod\limits_{v\in V}\mathsf{denom}(\lambda_{(v,\sigma(v))})$.
This value is easily estimated by using the highest possible denominators available in $\mathcal{G}$, and its representation is polynomial in the size of $\mathcal{G}$.
Biased sum of offsets.
We modify sharp games that are not improving. This can be done by redefining the function $\mathsf{offset}$ as follows:
$\mathsf{offset}^{\prime}(\mathsf{val},(v,v^{\prime}))=\alpha_{(v,v^{\prime})}\cdot\mathsf{offset}(\mathsf{val},(v,v^{\prime}))$
Such $\mathsf{offset}^{\prime}$ is defined for every edge $e=(v,v^{\prime})$, where each $\alpha_{e}$ is a positive number, which we call the offset factor. Based on this change, we re-define all offset definitions and the objective function with a primed version that uses these $\mathsf{offset}^{\prime}$ factors, and thus also what an improving DPG is.
Theorem 10.
Let $\mathcal{G}=(V_{\min},V_{\max},E,w,\lambda)$ be an improving DPG for a given set of positive numbers $\{\alpha_{e}>0\mid e\in E\}$, $\sigma$ a strategy for both players, $\mathsf{val}$ an optimal solution returned at Line 5 of Algorithm 2 for the adjusted function $f_{\sigma}^{\prime}$, and let there be no local improvements of $\sigma$ for $\mathsf{val}$. Then there is a neighbouring solution $\mathsf{val}^{\prime\prime}$ to $\mathsf{val}$ such that there is a better strategy $\sigma^{\prime}$ that satisfies $f_{\sigma^{\prime}}(\mathsf{val}^{\prime\prime})<f_{\sigma}(\mathsf{val})$.
This strategy $\sigma^{\prime}$ can be selected in such a way that $(v,\sigma^{\prime}(v))\in E_{\mathsf{val}}^{\sigma}$ holds for all $v\in V$ for a given set $E_{\mathsf{val}}^{\sigma}$.
The proof of Theorem 6 can be repeated verbatim with the new offset definition.
Theorem 11.
If each offset factor $\alpha_{e}$ is selected independently uniformly at random from a bounded interval of positive numbers888or from any other distribution over positive numbers that has 0 weight for all individual points, then a sharp DPG $\mathcal{G}=(V_{\min},V_{\max},E,w,\lambda)$ is almost surely improving for a sampled set of positive numbers $\{\alpha_{e}>0\mid e\in E\}$.
Proof 5.6.
We show the claim that, for two arbitrary but fixed (before drawing the $\alpha_{e}$) different solutions $\mathsf{val}_{1}$ and $\mathsf{val}_{2}$ that satisfy the inequations in $H$ and any joint strategy $\sigma$, $f_{\sigma}(\mathsf{val}_{1})\neq f_{\sigma}(\mathsf{val}_{2})$ holds almost surely.
We first assume for contradiction that $\mathsf{offset}(\mathsf{val}_{1},(v,\sigma(v)))=\mathsf{offset}(\mathsf{val}_{2},(v,\sigma(v)))$ holds for all $v\in V$. We pick a vertex $v$ such that $|\mathsf{val}_{1}(v)-\mathsf{val}_{2}(v)|>0$ is maximal; such a vertex exists since $\mathsf{val}_{1}\neq\mathsf{val}_{2}$. For this $v$, $\mathsf{offset}(\mathsf{val}_{1},(v,\sigma(v)))=\mathsf{offset}(\mathsf{val}_{2},(v,\sigma(v)))$ entails $\mathsf{val}_{1}(v)-\mathsf{val}_{2}(v)=\lambda_{(v,\sigma(v))}(\mathsf{val}_{1}(\sigma(v))-\mathsf{val}_{2}(\sigma(v)))$. Using $\lambda_{e}\in[0,1)$ for $e=(v,\sigma(v))$, we get $|\mathsf{val}_{1}(\sigma(v))-\mathsf{val}_{2}(\sigma(v))|>|\mathsf{val}_{1}(v)-\mathsf{val}_{2}(v)|>0$, which contradicts the maximality of $v$.
We therefore have that $\mathsf{offset}(\mathsf{val}_{1},(v,\sigma(v)))\neq\mathsf{offset}(\mathsf{val}_{2},(v,\sigma(v)))$ holds for some $v\in V$. We fix such a vertex $v$.
As the $\alpha_{e}$ are drawn independently, we can assume w.l.o.g. that $\alpha_{(v,\sigma(v))}$ is drawn last. There is at most one value $\alpha_{(v,\sigma(v))}^{\prime}$ for which the following condition is not satisfied:
$\sum_{v\in V}\mathsf{offset}^{\prime}(\mathsf{val}_{1},(v,\sigma(v)))\neq\sum_{v\in V}\mathsf{offset}^{\prime}(\mathsf{val}_{2},(v,\sigma(v))).$
It therefore holds almost surely that $f_{\sigma}(\mathsf{val}_{1})\neq f_{\sigma}(\mathsf{val}_{2})$.
As $\mathcal{G}$ is sharp, this implies for all joint strategies $\sigma$ and all neighbouring bases $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$ that define solutions $\mathsf{val}_{1}$ and $\mathsf{val}_{2}$, respectively, that $\mathsf{val}_{1}\neq\mathsf{val}_{2}$ (by the sharpness of $\mathcal{G}$), which we showed implies that $f_{\sigma}(\mathsf{val}_{1})\neq f_{\sigma}(\mathsf{val}_{2})$ holds almost surely. That is, it implies that $\mathcal{G}$ is improving for the sampled set of positive numbers.
Thus, we can almost surely obtain sharpness by adding small noise to the weights, and almost surely make games improving by considering the offsets of the individual edges with a randomly chosen positive weight. This guarantees cheap progress for the case where there are no local improvements.
@@ Acknowledgements
This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101032464. It was supported by the EPSRC through the projects EP/X017796/1 (Below the Branches of Universal Trees) and EP/X03688X/1 (TRUSTED: SecuriTy SummaRies for SecUre SofTwarE Development). We also thank Peter Austin from the University of Liverpool for providing an integration of the LP solver.

### v5

@@ 5.3. Games with Efficient Objective Improvement
The second condition, which allows us to identify better strategies efficiently, is to assume that, for every strategy $\sigma$ for both players, if a solution $\mathsf{val}$ defined by a basis is not optimal for $f_{\sigma}$ under the constraints $H$, then there is a single basis change that improves it. We call such games improving and show that all sharp games are improving.
@@
(Improving Game) A game $\mathcal{G}$ is called $\sigma$-improving for an objective function $f_{\sigma}$ defined by a joined strategy $\sigma$ and a non-optimal solution $\mathsf{val}$ defined by a basis for $f_{\sigma}$ if, and only if, there is a single basis change that leads to a better solution $\mathsf{val}^{\prime}$ (having a smaller value of $f_{\sigma}(\mathsf{val}^{\prime})<f_{\sigma}(\mathsf{val})$) defined by the new basis.
@@
Theorem 6.
Every sharp game $\mathcal{G}$ is improving.
Proof 5.3.
As our definition of sharp games follows simply says that every basic solution is nondegenerate, every step of a simplex algorithm (outside of optimal solutions) will provide a strict improvement [BT97, Theorem 3.3], so that $\mathcal{G}$ is $\sigma$-improving for every joined strategy $\sigma$ and all solutions $\mathsf{val}$ defined by bases that are not optimal w.r.t. $f_{\sigma}$.
@@
Let $\mathcal{G}=(V_{\min},V_{\max},E,w,\lambda)$ be an improving DPG, $\sigma$ a strategy for both players that is not co-optimal, $\mathsf{val}$ an optimal solution returned at Line 5 of Algorithm 2 for $f_{\sigma}$. Then there is (at least) one neighbouring solution $\mathsf{val}^{\prime\prime}$ to $\mathsf{val}$ such that there is a better strategy $\sigma^{\prime}$ that satisfies $f_{\sigma^{\prime}}(\mathsf{val}^{\prime\prime})<f_{\sigma}(\mathsf{val})$.
This strategy $\sigma^{\prime}$ can be selected in such a way that $(v,\sigma^{\prime}(v))\in E_{\mathsf{val}}^{\sigma}$ holds for all $v\in V$ for the given set $E_{\mathsf{val}}^{\sigma}$.
Proof 5.4.
@@
5.4. Making Games Sharp (and Thus Improving)
Naturally, not every game is sharp. In this subsection, we discuss how to almost surely make games sharp by adding sufficiently small random noise to the edge weights. Note that these are ‘global’ adjustments of the game that only need to be applied once, as it is the game that becomes sharp.
We first create notation for expressing how much we can change edge weights by adding small noise, such that joint co-optimal strategies of the resulting game are joint co-optimal strategies in the original game. To this end, we define the gap of a game.
@@ 5.4. Making Games Sharp (and Thus Improving)
This defines a run $\rho^{\prime}=d_{1}d_{2}d_{3}\ldots$ (with $d_{1}=e$), again in the form of a “lasso path” that consists of an initial path $d_{1},\ldots,d_{k^{\prime}}$, followed by an infinitely often repeated cycle $d_{1}^{\prime},\ldots,d_{\ell^{\prime}}^{\prime}$, where every edge on the path $d_{1},\ldots,d_{k^{\prime}},d_{1}^{\prime},\ldots,d_{\ell^{\prime}}^{\prime}$ occurs only once.
@@
This value is easily estimated by using the highest possible denominators available in $\mathcal{G}$—maximising $\mathsf{denom}(\lambda_{e})\cdot\mathsf{denom}(w_{e})$ over all edges $e\in E$ and maximising $\mathsf{denom}(\lambda_{(v,\sigma(v))})^{3}\cdot\mathsf{denom}(w_{(v,\sigma(v))})$ individually for every vertex $v\in V$. Its representation is polynomial in the size of $\mathcal{G}$.
Thus, we can almost surely obtain sharpness by adding small noise to the weights, and the resulting sharp games are also improving by Theorem 6. This guarantees cheap progress for the case where there are no local improvements.
@@ Acknowledgements
This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101032464. It was supported by the EPSRC through the projects EP/X017796/1 (Below the Branches of Universal Trees) and EP/X03688X/1 (TRUSTED: SecuriTy SummaRies for SecUre SofTwarE Development). We also thank Peter Austin from the University of Liverpool for providing an integration of the LP solver. We thank an anonymous reviewer for pointing out that all sharp games are improving, which simplifies our methods compared to the conference version and thus improves the results.